OpenStudy (anonymous):

What does it mean for a number to be the "solution" to an equation? Give an example of a variable equation and its solution. Explain in complete sentences how you know that number is the solution.

OpenStudy (anonymous):

@Valkarie70

OpenStudy (anonymous):

a number is a "solution to an equation" means if you replace the variable by that number, the equation is true. a simple example would be \(x+2=7\) the solution is \(5\) because \(5+2=7\) is correct if you replace \(x\) by any number other than 5 it would not be correct

OpenStudy (anonymous):

3x+20-5=21 x=2

OpenStudy (anonymous):

exactly what satellite73 said is correct